# Boosting Adversarial Transferability with Spatial Adversarial Alignment

### v1

@@ 4.1 Experimental Setup
Attack Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT Avg. ASR (%)
MI 57.7 99.9 58.1 54.2 55.1 39.0 26.3 15.1 38.0 35.7 42.1
MI-SAA 84.1 99.6 74.7 80.3 81.8 65.7 53.1 34.8 52.9 62.5 65.5
NI 58.9 100.0 63.2 59.3 61.4 40.0 27.3 15.2 41.8 38.1 45.0
NI-SAA 86.1 99.9 76.3 82.2 83.7 67.6 53.2 35.6 55.7 64.8 67.2
GI 57.3 100.0 62.3 60.5 60.5 40.7 29.5 17.3 40.7 39.6 45.4
GI-SAA 86.5 99.7 78.8 83.9 84.8 70.4 58.8 39.6 55.8 66.3 69.4
DI 44.1 95.8 41.7 56.1 44.2 26.1 16.5 7.7 36.7 35.1 34.2
DI-SAA 74.6 94.4 61.1 81.1 73.2 53.1 34.9 19.9 50.8 63.6 56.9
TI 38.5 99.9 37.8 33.9 36.1 24.2 16.8 7.9 29.0 21.1 27.3
TI-SAA 59.7 94.9 35.2 50.6 54.5 40.6 29.2 13.6 29.3 33.6 38.5
SSA 75.8 99.9 78.6 76.0 77.8 57.0 35.3 24.1 55.0 50.5 58.9
SSA-SAA 91.5 99.5 77.8 85.7 88.4 74.9 54.4 36.9 57.1 66.0 70.3
DI-MI 65.5 97.0 65.0 74.7 65.7 49.1 34.5 22.1 54.9 57.9 54.4
DI-MI-SAA 91.9 98.7 84.9 94.1 90.5 78.3 64.9 46.2 76.1 86.8 79.3
TI-MI 61.4 99.9 60.5 60.9 60.9 44.3 35.5 19.6 42.3 41.8 47.5
TI-MI-SAA 84.8 99.3 71.9 79.0 81.8 69.1 58.9 38.7 52.8 62.4 66.6
SSA-MI 89.6 99.9 92.2 89.5 91.0 77.6 57.8 49.4 76.4 76.3 77.8
SSA-MI-SAA 96.3 99.8 95.6 96.5 97.2 91.5 81.3 68.2 80.4 88.4 88.4
SSA-DI-TI-MI 93.5 98.5 92.3 95.0 93.7 85.5 78.9 66.5 87.1 89.8 86.9
SSA-DI-TI-MI-SAA 97.5 98.8 93.8 97.6 96.7 94.3 90.6 78.0 84.4 94.7 92.0
@@
Attack Target Model Avg. ASR (%)
@@
4.2 Performance Comparison
Performance comparison with alignment methods. We first compare with existing alignment methods [26], where adversarial examples are generated based on MI [9]. Table 2 illustrates the performance difference between MA and SAA in terms of adversarial transferability, with SAA demonstrating a significant advantage over MA. For instance, when the surrogate model is Res50, and the witness model is also Res50, SAA achieves a 16.6% improvement in average ASR over the original surrogate model, compared to a modest 3.6% improvement with MA. This highlights that SAA, even without introducing additional information, enhances adversarial transferability through the alignment of adversarial features. Furthermore, when the witness models are DN121, ViT-B, and Swin-B, SAA outperforms MA by 11.9%, 10.4%, and 13.1%, respectively. In addition to the remarkable adversarial transferability that SAA provides, we make two other key observations: (i) MA only considers global features, which makes it difficult to align features between models with large differences, which may lead to a decrease in transferability. When the surrogate model is DN121 and the witness model is Swin-B, the ASR of Vit-B, Swin-B, PVT-v2, and MobViT is not as good as the origin DN121, which shows that relying solely on global features for alignment is not enough, and can only achieve poor transferability. (ii) SAA provides strong cross-architecture transferability. When the surrogate model is Res50, and the witness models are Res50, DN121, ViT-B, and Swin-B, the transferability of SAA on ViTs is improved by 39.06%, 31.29%, 25.51%, and 37.74% respectively compared with MA itself, and it also has high transferability between CNNs.
@@ 4.2 Performance Comparison
Module Target Model Avg. ASR (%)
@@
Performance comparison with transfer attacks. Aligned surrogate models by SAA have great potential for adversarial transferability, so existing transfer attacks such as advanced optimization and data augmentation can further improve transferability. Here, we choose Res50 as the surrogate model and ViT-B as the witness model, and superimpose them with MI [9], NI [22], GI [37], DI [45], TI [10] and SSA [25] to evaluate the transferability, as shown in Table 3. Taking GI and SSA as examples, the transferability of the model after SAA is improved by 24.0% and 11.4% respectively compared with the origin surrogate model, which is a very significant improvement. When multiple attacks are integrated, such as SSA-DI-TI-MI, SAA further enhances the transferability by 5.1%, achieving an impressive 92.0% ASR, which closely approaches the performance of white-box attacks. This indicates that SAA substantially narrows the performance gap between black-box and white-box attacks, thereby facilitating a more comprehensive evaluation of the adversarial robustness of existing models.
Performance comparison with attacks on ViTs. To further explore the cross-architecture transferability, we evaluate adversarial attacks on ViTs, including SGM [44], PatchOut [43], PNA [43], and TGR [49]. Here, we choose the surrogate model as ViT-B and the witness model as Res50. In PatchOut, SAA first improves the transferability between ViTs, for example, the ASR from ViT-B to Swin-B is improved by 42.6%. Secondly, SAA greatly improves the transferability from ViT to CNN, for example, the ASR is improved by 45.0% and 42.0% on Res50 and DN121 respectively. On SGM, PNA, and TGR, SAA also achieves stronger cross-architecture transferability without modifying the forward and backward propagation of the model.
4.3 Ablation Studies
We select ResNet-50 as the surrogate model and ViT-B as the witness model for ablation studies.
Alignment Module. In spatial-aware alignment, ‘global’ represents $\mathcal{L}_{global}$ (Equation 1), while ‘local’ represents $\mathcal{L}_{local}$ (Equation 2). Similarly, ‘adversarial’ represents $\mathcal{L}_{AA}$ (Eqution 5) of adversarial-aware alignment. As shown in Table 5, when global features are introduced into the alignment, the transferability of the aligned surrogate model will increase by 3.6%. Based on $\mathcal{L}_{global}$, when only local features are introduced, the overall transferability is improved by 4.3% due to better alignment of features of different architectures in local regions, especially by 8.7% on MobViT. Based on $\mathcal{L}_{global}$, when only adversarial features are introduced, the transferability is greatly improved, reaching 12.6% ASR, and the improvement is significant on ViTs. Finally, by integrating all features, the aligned surrogate model achieves state-of-the-art transferability. Through the above experiments, we illustrate the effectiveness of spatial-aware and adversarial-aware alignment.

### v2

@@ 3.2 Performance Comparison
Attack Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT
Avg.
ASR (%)
MI 57.7 99.9 58.1 54.2 55.1 39.0 9.4 33.0 38.0 35.7 42.2
MI-SAA 84.1 99.6 74.7 80.3 81.8 65.7 24.3 48.7 52.9 62.5 63.9
NI 58.9 100.0 63.2 59.3 61.4 40.0 9.6 37.4 41.8 38.1 45.5
NI-SAA 86.1 99.9 76.3 82.2 83.7 67.6 24.0 50.6 55.7 64.8 65.7
GI 57.3 100.0 62.3 60.5 60.5 40.7 12.5 36.8 40.7 39.6 45.7
GI-SAA 86.5 99.7 78.8 83.9 84.8 70.4 27.6 52.7 55.8 66.3 67.4
DI 44.1 95.8 41.7 56.1 44.2 26.1 5.6 30.9 36.7 35.1 35.6
DI-SAA 74.6 94.4 61.1 81.1 73.2 53.1 10.6 44.0 50.8 63.6 56.9
TI 38.5 99.9 37.8 33.9 36.1 24.2 5.4 21.0 29.0 21.1 27.4
TI-SAA 59.7 94.9 35.2 50.6 54.5 40.6 9.5 22.8 29.3 33.6 37.3
SSA 75.8 99.9 78.6 76.0 77.8 57.0 16.5 48.5 55.0 50.5 59.5
SSA-SAA 91.5 99.5 77.8 85.7 88.4 74.9 23.4 46.7 57.1 66.0 67.9
DI-MI 65.5 97.0 65.0 74.7 65.7 49.1 16.4 49.0 54.9 57.9 55.4
DI-MI-SAA 91.9 98.7 84.9 94.1 90.5 78.3 34.1 69.8 76.1 86.8 78.5
TI-MI 61.4 99.9 60.5 60.9 60.9 44.3 15.2 37.4 42.3 41.8 47.2
TI-MI-SAA 84.8 99.3 71.9 79.0 81.8 69.1 27.0 45.0 52.8 62.4 63.8
SSA-MI 89.6 99.9 92.2 89.5 91.0 77.6 39.2 74.4 76.4 76.3 78.5
SSA-MI-SAA 96.3 99.8 95.6 96.5 97.2 91.5 46.3 74.1 80.4 88.4 85.1
SSA-DI-TI-MI 93.5 98.5 92.3 95.0 93.7 85.5 55.9 83.2 87.1 89.8 86.2
SSA-DI-TI-MI-SAA 97.5 98.8 93.8 97.6 96.7 94.3 53.6 81.6 84.4 94.7 88.2
@@
Attack Target Model
Avg.
ASR (%)
@@
Table 5: SAA improves the adversarial transferability of ensemble attacks.
Target Model
CNNs ViTs
Surrogate Attack Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT Avg. ASR (%)
Res50 n/a 57.7 99.9 58.1 54.2 55.1 39.0 9.4 33.0 38.0 35.7 42.2
ViT-B n/a 48.6 41.0 31.9 58.1 50.8 44.4 100.0 58.1 44.7 47.4 47.2
Res50+ViT-B n/a 61.3 100.0 63.0 60.8 61.2 45.0 97.6 50.4 51.3 48.1 55.1
[…]
Res50+DN121+ViT-B+Swin-B n/a 65.4 100.0 78.0 79.1 99.8 62.0 93.3 100.0 77.4 68.9 71.8
[…]
Res50(ViT-B)+Res50(Swin-B)
SAA 93.5 100.0 92.9 92.1 95.1 82.3 39.6 71.8 80.4 81.7 87.2
Table 6: SAA improves adversarial transferability against adversarial defenses.
Attack HGD R&P NIPS-r3 JPEG FD RS Bit-Red NRP Diffpure Inc-v3ens3 Inc-v3ens4 IncRes-v2ens Avg. ASR (%)
MI 42.5 21.9 25.3 33.9 42.4 23.6 29.3 6.7 13.8 33.3 31.3 23.3 27.3
MI-SAA 73.3 57.8 60.4 69.9 65.6 39.6 42.3 12.0 22.4 68.9 65.5 57.9 53.0
SSA-DI-TI-MI 93.7 89.6 90.2 91.9 89.7 82.3 81.7 14.8 71.1 92.5 91.1 89.8 81.5
SSA-DI-TI-MI-SAA 96.0 93.2 94.8 95.1 94.0 89.8 85.5 20.2 78.8 95.7 94.5 93.3 85.9
Performance comparison with transfer attacks. Aligned surrogate models by SAA have great potential for adversarial transferability, so existing transfer attacks such as advanced optimization and data augmentation can further improve transferability. Here, we choose Res50 as the surrogate model and ViT-B as the witness model, and superimpose them with MI mi , NI sini , GI WANG2024124757 , DI xie2019improving , TI dong2019evading and SSA ssa2022 to evaluate the transferability, as shown in Table 3. Taking GI and SSA as examples, the transferability of the model after SAA is improved by 21.7% and 8.4% respectively, compared with the origin surrogate model, which is a very significant improvement. When multiple attacks are integrated, such as SSA-DI-TI-MI, SAA further enhances the transferability by 2.0%, achieving an impressive 88.2% ASR, which closes white-box attacks’ performance. This indicates that SAA substantially narrows the performance gap between black-box and white-box attacks, thereby facilitating a more comprehensive evaluation of the adversarial robustness of existing models.
Performance comparison with attacks on ViTs. To further explore the cross-architecture transferability, we evaluate adversarial attacks on ViTs, including SGM sgm , PatchOut pna , PNA pna , and TGR tgr . Here, we choose the surrogate model as ViT-B and the witness model as Res50. In PatchOut, SAA first improves the transferability between ViTs, for example, the ASR from ViT-B to Swin-B is improved by 42.6%. Secondly, SAA greatly improves the transferability from ViT to CNNs, for example, the ASR is improved by 45.0% and 42.0% on Res50 and DN121 respectively. On SGM, PNA, and TGR, SAA also achieves stronger cross-architecture transferability without modifying the forward and backward propagation of the model.
Summary. Although we conduct 16 combinations and evaluate on 6 CNNs and 4 ViTs, showing that SAA’s improvement is not limited to specific choices of witness models. Based on the above experiments, we further summarize the empirical guidance for the selection of surrogate and witness models: i) Self-alignment (surrogate and witness models are consistent) usually only provides minor improvements (it is difficult to learn unique structural features), and alignment between different models improves more significantly. ii) When the surrogate model is a ViT-like model and the witness model is a CNN-like model, better cross-architecture transferability can be obtained. This is mainly because ViT has stronger performance, while CNN further provides unique structural features. It should be noted that the target models are unknown in the transfer attack, so no information of the target models can be used for alignment.
3.3 Ensemble Attacks
The ensemble of multiple surrogate models can further improve adversarial transferability mi , so we want to explore whether SAA can further improve transferability. In the setting of ensemble attacks, we follow the setting of MI mi and use the logits of multiple surrogate models for integration. Here we choose multiple groups of settings, such as Res50 (ViT-B) means that the surrogate model is Res50 and the witness model is ViT-B. Res50+ViT-B means the logits of these two models are integrated, and so on. As shown in Table 5, SAA further improves the performance of ensemble attacks. First, the alignment-based method is not as good as the direct integration strategy in white-box performance, but it has significant improvement in black-box transferability. For example, on (Res50+ViT-B), MA and SAA are 9.1% and 20.8% higher than the direct ensemble ASR, respectively. Then, MA directly constrains global features and does not take into account the huge differences in features between different structures, which may lead to degradation of transferability, such as on Swin-B. Finally, SAA has achieved state-of-the-art transferability and further enhance ensemble attacks because it considers spatial and adversarial features from a structural perspective.
3.4 Adversarial Defenses
To illustrate the effect of SAA in the face of adversarial defenses, we choose Inc-v3ens3 as the target model and evaluate its performance on 12 adversarial defenses, including HGD DBLP:conf/cvpr/LiaoLDPH018 , R&P xie2017mitigating , NIPS-r3222https://github.com/anlthms/nips-2017/tree/master/mmd, JPEG guo2017countering , FD liu2019feature , RS DBLP:conf/icml/CohenRK19 , Bit-Red DBLP:conf/ndss/Xu0Q18 , NRP DBLP:conf/cvpr/NaseerKHKP20 , Diffpure nie2022DiffPure , Inc-v3ens3 Florian2018Ensemble , Inc-v3ens4 Florian2018Ensemble , and IncRes-v2ens Florian2018Ensemble . Here we use the ensemble attack setting, and the surrogate models are Res50(Res50)+Res50(DN121)+Res50(ViT-B)+Res50(Swin-B). As shown in Table 6, although adversarial defenses weaken transferability, SAA still achieves a significant improvement in adversarial transferability compared to the origin surrogate models.
3.5 Ablation Studies
We select ResNet-50 as the surrogate model and ViT-B as the witness model for ablation studies (see Appendix C for more details, including training epochs, distance metric, GPU memory and computing cost, self-adversarial strategy).
Alignment Module. In spatial-aware alignment, ‘global’ represents $\mathcal{L}_{global}$ (Equation 1), while ‘local’ represents $\mathcal{L}_{local}$ (Equation 2). Similarly, ‘adversarial’ represents $\mathcal{L}_{AA}$ (Eqution 5) of adversarial-aware alignment. As shown in Table 7, when global features are introduced into the alignment, the transferability of the aligned surrogate model will increase by 3.6%. Based on $\mathcal{L}_{global}$, when only local features are introduced, the overall transferability is improved by 4.3% due to better alignment of features of different architectures in local regions, especially by 8.7% on MobViT. Since intermediate layer features between models are diverse and different, we choose last layer features that summarize the spatial information, which is a simple and effective strategy. Based on $\mathcal{L}_{global}$, when only adversarial features are introduced, the transferability is greatly improved, reaching 12.6% ASR, and the improvement is significant on ViTs. Finally, by integrating all features, the aligned surrogate model achieves state-of-the-art transferability.
@@ 3.5 Ablation Studies
Module Target Model
Avg.
ASR (%)
@@
4 Conclusions
In this study, we introduce a novel technique called Spatial Adversarial Alignment (SAA), which incorporates an alignment loss function and utilizes a witness model to fine-tune a surrogate model by focusing on both spatial-aware and adversarial-aware alignments. Through experimental analysis, we show that leveraging these spatial and adversarial features for model alignment significantly enhances the adversarial transferability of surrogate models, with a particularly pronounced improvement in their cross-architecture capabilities. SAA not only integrates seamlessly with existing transfer attack strategies but also further amplifies adversarial transferability, thereby contributing to a more complete evaluation of the adversarial robustness of DNNs.
Boarder Impacts. Adversarial examples by SAA exhibit enhanced adversarial transferability, especially in cross-architecture capabilities. This poses a huge threat to the deployment of real-world applications. Simultaneously, it is also conducive to better evaluating their adversarial robustness.
Limitations. The huge gaps between network architectures limit transferability. While SAA aligns unique features under both spatial and adversarial conditions to mitigate these gaps, it does not fully resolve them. Then, SAA also introduces some computing costs during the alignment but no extra costs during attacking. In addition, a theoretical analysis is lacking. We have made significant progress, but there is still much to be done in addressing this issue.
Acknowledgments. This work was supported by National Natural Science Foundation of China (No.62576109, 62072112).
Appendix A Related Work
A.1 Transfer attacks on CNNs
A.2 Transfer attacks on ViTs
Algorithm 1 Loss $\mathcal{L}_{SAA}$
1:Input $x$, surrogate model $f_{\theta_{s}}$, witness model $f_{\theta_{w}}$
[…]
3:#Preliminaries
4:if $f_{\theta}$ is CNN then
5:  Compute global logits:
6:    $f_{\theta}(x)=\text{Classifier}(\text{GlobalAvgPooling}(Backbone(x)))$
7:  Obtain local regions collection $z$:
8:    $z_{\theta}(x)=\text{Classifier}(Backbone(x))$
9:else if $f$ is ViT then
10:  $z^{\prime}=\text{TransformerEncoder}(x)$
11:  Compute global logits:
12:    $f_{\theta}(x)=\text{Classifier}(z^{\prime}[0])$
13:  Obtain local regions collection $z$:
14:    $z_{\theta}(x)=\text{Classifier}(z^{\prime}[1:])$
15:end if
16:#Step 1: Compute spatial-aware alignment loss
17:Obtain global logits $f_{\theta_{s}}(x)$ and $f_{\theta_{w}}(x)$
18:Compute $\mathcal{L}_{global}(x;\theta_{s})=D_{\mathrm{KL}}(f_{\theta_{s}}(x),f_{\theta_{w}}(x))$
19:for each local region $q$ do
20:  Obtain local logits $z_{\theta_{s}}^{[q]}$ and $z_{\theta_{w}}^{[q]}$
21:  Derive pseudo-label $\hat{y}^{[q]}_{\theta_{w}}=\arg\max(z_{\theta_{w}}^{[q]}(x))$
[…]
23:end for
24:Calculate total local alignment loss: $\mathcal{L}_{local}(x;\theta_{s})=\frac{1}{HW}\sum^{HW}_{q=1}D_{\mathrm{CE}}(z_{\theta_{s}}^{[q]}(x),\arg\max(z_{\theta_{w}}^{[q]}(x)))$
25:Calculate spatial-aware alignment loss: $\mathcal{L}_{SA}(x;\theta_{s})=\mathcal{L}_{global}(x;\theta_{s})+\gamma\cdot\mathcal{L}_{local}(x;\theta_{s})$
26:#Step 2: Compute adversarial-aware alignment loss
27:Initialize adversarial example $x_{\text{adv}}^{(0)}=x$
28:for each iteration $t$ do
29:  $x_{adv}^{(t+1)}=\Pi_{x,\epsilon}\left(x_{\text{adv}}^{(t)}+\alpha\cdot\text{sign}\left(\nabla_{x}D_{\text{KL}}\left(f_{\theta_{s}}(x_{\text{adv}}^{(t)}),f_{\theta_{w}}(x)\right)\right)\right)$
30:end for
31:Obtain final adversarial example $x_{\text{adv}}=x_{\text{adv}}^{(T)}$
32:Calculate adversarial-aware alignment loss: $\mathcal{L_{AA}}(x_{adv};\theta_{s})=\mathcal{L}_{global}(x_{adv};\theta_{s})+\omega\cdot\mathcal{L}_{local}(x_{adv};\theta_{s})$
33:#Step 3: Calculate spatial adversarial alignment loss
[…]
35:Return $\mathcal{L}_{SAA}$
Appendix B Loss Calculation
We introduce the calculation of loss in detail, as shown in Algorithm 1. For the specific implementation, please refer to the code provided in Supplementary Material.
Table 8: Ablation study on distance metrics of global features.
Loss Target Model Avg. ASR (%)
CNNs ViTs
Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT
n/a 57.7 99.9 58.1 54.2 55.1 39.0 9.4 33.0 38.0 35.7 42.2
KL 84.1 99.6 74.7 80.3 81.8 65.7 24.3 48.7 52.9 62.5 63.9
TV 62.1 97.8 56.2 67.5 63.9 47.6 9.4 32.2 37.2 44.1 46.7
JS 83.3 99.1 70.0 78.2 80.9 63.5 22.8 46.4 49.3 58.5 61.4
NCE 78.8 99.3 69.1 77.0 76.6 60.3 21.7 46.2 48.2 55.8 59.3
Table 9: Ablation study on different supervision in self-adversarial strategy.
Source Witness Attack Res18 Res50 Res101 VGG19 DN121 IncV3 ViT-B Swin-B PVT-v2 MobViT Avg. ASR (%)
Res50 ViT-B MA 74.2 99.2 63.5 69.3 72.8 51.5 18.5 41.5 42.7 47.4 53.5
SAA ($f_{\theta_{w}}(x)$) 84.1 99.6 74.7 80.3 81.8 65.7 24.3 48.7 52.9 62.5 63.9
SAA ($f_{\theta_{s}}(x)$) 81.1 96.6 68.7 77.5 79.5 60.6 19.4 45.9 46.2 57.6 59.6
Table 10: Ablation study on label supervision in local alignment.
Source Witness Attack Res18 Res50 Res101 VGG19 DN121 IncV3 ViT-B Swin-B PVT_v2 MobViT Avg. ASR (%)
Res50 ViT-B SAA-MI (Ours) 84.1 99.6 74.7 80.3 81.8 65.7 24.3 48.7 52.9 62.5 63.9
SAA-MI (soft) 81.9 96.4 65.9 74.5 77.7 60.8 18.8 44.4 43.1 53.0 57.8
Table 11: SAA improves the adversarial transferability of feature-based transfer attacks.
Attack Target Model Avg. ASR (%)
CNNs ViTs
Res18 Res50 Res101 VGG19 DN121 Inc-v3 ViT-B Swin-B PVT-v2 MobViT
FDA 55.2 31.6 36.8 52.0 43.6 27.1 2.7 21.0 24.5 27.8 32.3
FDA-SAA 82.9 79.0 72.5 79.1 77.1 60.1 14.4 39.5 43.4 58.4 58.6
FDA-MI 75.0 46.1 60.6 74.0 65.3 46.0 8.4 27.3 29.4 47.0 48.1
FDA-MI-SAA 89.7 88.2 81.7 86.9 85.3 73.1 21.9 41.7 44.7 68.9 66.0
Appendix C Ablation Study
We select ResNet, DenseNet, and Swin, all of which have won Best Paper Awards, and ViT, which is a pioneering work. These models are representative of both CNNs and ViTs and their weights are all from timm, ensuring the generalization of our conclusions. We conduct 16 combinations of experiments and evaluate on 6 CNNs and 4 ViTs, showing that SAA’s improvement is not limited to specific choices of witness models.
Figure 3: Ablation study on training epochs.